# Video Event Reasoning and Prediction by Fusing World Knowledge from LLMs with Vision Foundation Models

### v1

@@ 1 Introduction
The proliferation of video data has positioned it as a primary medium for information exchange and environmental perception, driving a significant evolution in computer vision research (9; 55). Historically, the field has achieved remarkable success in discriminative tasks—moving from foundational image recognition (6) to complex action and gesture recognition in videos (1). This progress is mirrored in specialized domains such as sign language translation, where hierarchical models have demonstrated powerful recognition capabilities (16; 18), and in fine-grained sensing technologies like WiFi-based gesture and activity recognition, which have become increasingly robust and resistant to interference (14; 25; 24; 62). However, this paradigm of "recognition" primarily addresses the question of "what" is happening in a visual scene. The community is now shifting towards a more profound challenge: moving from mere perception to genuine cognition (60). This involves enabling machines to reason about "why" an event is occurring and to predict "what" might happen next—a task that requires a level of understanding far beyond statistical pattern matching.
@@
The primary obstacle to achieving this cognitive leap is the "knowledge gap." Current models, despite their architectural sophistication, often operate within a closed world, lacking the vast repository of commonsense, physical intuition, and social knowledge that humans effortlessly apply. For instance, while a model might accurately classify a video clip as "a person picking up an egg and flour," it typically fails to infer the underlying intent, such as "the person is about to bake a cake." This limitation becomes particularly salient when considering the demands of next-generation applications. Advanced human-machine systems (62), affective computing that aims to suppress label noise for reliability (63), and even emerging fields like psychological understanding via language models (23) all necessitate a deeper, causal understanding of events. Similarly, while we can now capture nuanced physiological data through commodity devices (65; 50), interpreting this data in the context of complex human activities requires advanced reasoning. The core challenge, therefore, is to imbue visual models with this external world knowledge.
@@
To bridge this gap, we propose a novel framework that synergistically fuses the capabilities of two of the most powerful paradigms in modern AI: Vision Foundation Models (VFMs) and Large Language Models (LLMs). Our approach leverages VFMs, such as those built upon the principles of Vision Transformers (6) and trained with multimodal supervision (41; 52), to serve as the "eyes" of our system. These models are adept at extracting rich, spatiotemporal features and providing a detailed, pixel-level perception of the visual world. Concurrently, we employ a pre-trained LLM, such as LLaMA (47), as the "brain"—a reasoning core endowed with a vast reservoir of world knowledge, causal relationships, and abstract concepts, as demonstrated in seminal works like GPT-4 (4). The crux of our methodology lies in a meticulously designed fusion mechanism that translates the continuous, unstructured visual evidence from the VFM into a discrete, language-like format that the LLM can process and reason upon. This approach is inspired by the success of pioneering vision-language architectures like Flamingo (2) and BLIP-2 (33), but is explicitly tailored for complex event-level inference and prediction, moving beyond simple captioning or direct question-answering.
@@
This work makes several significant contributions to the advancement of cognitive video understanding. First and foremost, we introduce a novel and effective framework that, for the first time, cohesively integrates a state-of-the-art Vision Foundation Model with a large-scale LLM to perform high-level event reasoning and prediction, moving decisively beyond simple recognition tasks. Central to this framework is our second contribution: the design of a lightweight yet powerful cross-modal fusion module. This component acts as an efficient information bottleneck, aligning rich visual features with the LLM’s semantic space, which not only enables complex inference but also ensures that the model’s reasoning is firmly grounded in direct visual evidence. To validate our approach, our third contribution is a set of extensive experiments conducted on multiple challenging video reasoning benchmarks. The results demonstrate that our model significantly outperforms existing state-of-the-art methods and, notably, exhibits remarkable zero-shot capabilities in predicting future events, underscoring the profound benefits of transferring world knowledge from LLMs. We believe the principles developed herein also hold considerable promise for enhancing related multimodal tasks, such as visual dialog (19) and audio-visual event analysis (67).
@@ 2.1 Foundations of Visual Representation Learning
The journey towards meaningful video understanding begins with the extraction of powerful visual representations. Early successes were largely dominated by Convolutional Neural Networks (CNNs), which demonstrated exceptional capabilities in hierarchical feature extraction for images. However, the advent of the Transformer architecture, particularly the Vision Transformer (ViT) (6), marked a paradigm shift. By treating an image as a sequence of patches, ViTs enabled the application of self-attention mechanisms to capture global context, a feat that was challenging for CNNs with their limited receptive fields. This architectural innovation laid the groundwork for a new generation of foundation models. For video, this principle was extended to the temporal domain, leading to powerful video foundation models like InternVideo (52), which learn generalizable representations from massive datasets through a combination of generative and discriminative objectives. The development of such backbones is a field of research in itself, with ongoing efforts to improve efficiency and effectiveness, for example, through advanced model compression techniques like multi-objective convex quantization (7) or by designing specialized architectures for specific tasks like crowd counting (17).
@@ 2.2 From Recognition to Spatiotemporal Understanding
Building upon these powerful visual backbones, research has moved beyond simple classification towards a more nuanced understanding of spatiotemporal dynamics. This evolution is evident in tasks that require localizing events in both space and time. Video grounding, which aims to find a specific video segment corresponding to a textual query, is a prime example. Recent works like (22) have focused on developing efficient temporal filtering mechanisms to precisely identify these moments. The ambition has further expanded to generating structured summaries of long-form videos, such as creating distinct chapters, a task addressed by large-scale datasets and models like VidChapters-7M (10). This fine-grained temporal understanding is fundamental to our work, as reasoning about event causality and prediction necessitates a precise grasp of "when" things happen. A related task, text-to-video retrieval, further underscores the importance of fine-grained alignment, with recent benchmarks like Ground-A-Video (27) pushing the state of the art in accurately matching semantic queries to video content.
@@ 2.3 The Rise of Vision-Language Models (VLMs)
The true catalyst for advanced visual reasoning was the effective fusion of vision and language. The development of CLIP (41) demonstrated that a shared embedding space for images and text, learned via massive-scale contrastive pre-training, could enable remarkable zero-shot transfer capabilities. This breakthrough paved the way for a host of large-scale Vision-Language Models (VLMs). Early influential models like Flamingo (2) introduced gated cross-attention layers to inject visual features into a pre-trained and frozen language model, showcasing impressive few-shot learning. This "frozen LLM" paradigm was further explored in works like (48), which highlighted the potential of this frugal approach. Architectures like BLIP-2 (33) advanced this idea by introducing a lightweight "Q-Former" module to bridge the modality gap between a frozen image encoder and a frozen LLM, proving to be a highly effective and parameter-efficient strategy. The field continues to expand the scope of fusion, aiming to create omni-perception models like VALOR (5) and LanguageBind (68), which align not just vision and text, but also audio, depth, and thermal data into a unified semantic space. This trend towards multimodal fusion is not limited to mainstream sensors; innovative research has shown the potential of fusing commodity WiFi signals with vision for tasks like emotion recognition (15), illustrating a broader principle of synergistic sensing that our work builds upon.
@@ 2.4 Large Language Models for Video Understanding and Reasoning
The confluence of powerful VLMs and the demonstrated reasoning capabilities of LLMs (47; 4) has given rise to the current frontier: LLM-powered video understanding. A first wave of models, often framed as "video assistants," focused on enabling dialogue about video content. Models such as Video-LLaMA (64), Video-ChatGPT (38), and Chat-UniVi (29) demonstrated how a video encoder could be connected to an LLM to answer questions, generate descriptions, and hold conversations about a video’s content. LLaViDA (66) further explored enhancing this understanding via in-context learning.
Subsequently, research has pivoted towards enabling more complex and structured reasoning. SeViLA (59) introduced a self-chained question-answering approach, encouraging the model to break down problems into smaller, manageable steps. This aligns with broader trends in NLP, such as training models along explicit reasoning paths (13). Perhaps the most innovative approach is ViperGPT (45), which empowers an LLM to write and execute Python code that calls upon various vision APIs, effectively transforming the LLM into a cognitive orchestrator that can answer complex visual queries by composing modular tools. As the complexity of reasoning increases, so does the demand for handling longer contexts. Models like LaVi-L (61) and the memory-augmented Stammer (34) are specifically designed to address the challenges of long-form video understanding, which is crucial for tracing causal chains over extended periods.
This rapid progress has also necessitated a critical examination of model limitations, particularly the issue of "hallucination," where models generate factually incorrect or ungrounded text. Research like Woodpecker (58) is now focused on developing methods to detect and correct these hallucinations, a critical step towards building reliable systems. The ultimate ambition is to create unified, any-to-any multimodal models like Emu2 (44), NExT-GPT (54), and Google’s Gemini (12), which aim to seamlessly process and generate content across nearly all modalities. This includes extending reasoning into the third dimension, as explored by Chat-3D-v2 (56), and leveraging novel fusion architectures like AMAM’s modality-adaptive mind (30). The fusion principles explored in these works can even find analogues in other domains, such as the multi-view, multi-epoch architecture of SUTRA for speech processing (42), suggesting a universal trend in multimodal AI.
@@ 2.5 Event Prediction and World Models
The "prediction" component of our work is directly related to the long-standing challenge of video forecasting. Traditional approaches often focused on low-level prediction, such as generating future pixels. Diffusion models, as seen in MCVD (49), have recently shown great promise in generating high-fidelity future frames. However, our focus is on higher-level semantic prediction. This aligns with research in human trajectory prediction, where models like V-STF (28) learn to anticipate future movements by fusing social and temporal cues.
The most ambitious vision for prediction is embodied by the concept of "World Models." Pioneering work like DreamerV3 (20) has shown that an agent can learn a robust internal model of its environment’s dynamics, allowing it to "dream" or simulate future outcomes to plan its actions effectively. This represents a shift from reactive prediction to proactive simulation. The recent Genie model (46) takes this a step further, learning to generate entire interactive, playable 2D worlds from a single image. While our work does not build an explicit world model, it shares the same spirit: leveraging accumulated knowledge to make informed predictions about future states. The remarkable capabilities of generative models like VideoPoet (32) to synthesize coherent, dynamic video from text further suggest that these models are implicitly learning deep, predictive representations of the world.
@@ 2.6 Applications, Benchmarks, and Broader Context
The ultimate goal of video reasoning and prediction is to enable intelligent applications and systems. A major beneficiary is embodied AI and robotics. The paradigm has shifted from passive video analysis to training active agents that can perceive, reason, and act in the physical world. Landmark models like RT-2 (3) and the generalist Octo transformer (40) demonstrate that a single vision-language-action model can be trained to control robots for a variety of tasks. This requires not only understanding instructions but also organizing and planning complex actions, a challenge addressed by agents like LEO (37). The importance of external knowledge in these embodied tasks is highlighted by specialized benchmarks like OK-VILA (43).
Progress in this field is critically dependent on challenging and well-designed benchmarks. Datasets like CLEVRER (57) specifically target causal and physical reasoning, while Test of Time (39) focuses on evaluating temporal understanding. The Ego-Exo4D dataset (11) pushes the frontier by providing synchronized first-person and third-person views of the same event, demanding a more holistic, cross-view understanding. While much research relies on traditional visual data, parallel advancements in alternative sensing modalities are creating new opportunities. Technologies using commodity WiFi (14) and RFID (8) are now capable of fine-grained activity and even keystroke detection (26). These rich data streams, often generated in complex real-world settings like robotic vehicle perception (51), require the same, if not more, sophisticated reasoning models to be interpreted meaningfully. This is especially true in healthcare, where vision-based assessment of Parkinson’s tremors (53) or WiFi-based pulmonary function analysis (65) requires a deep understanding of subtle temporal patterns.
Finally, training these massive models on diverse, real-world data presents its own challenges, leading to research in areas like federated learning to handle decentralized data and heterogeneous networks, with frameworks like Finch enabling neural architecture search in such settings (36; 35). Our work is situated at the intersection of these advancements, aiming to leverage foundational models and advanced reasoning to create a system that not only understands video but can also anticipate its future, with broad implications for all these application domains.
@@ 3.1.1 Spatiotemporal Feature Extraction
We employ a pre-trained video foundation model, specifically a variant of the Vision Transformer (ViT) architecture adapted for video, such as InternVideo (52), as our primary spatiotemporal encoder. The input video $V\in\mathbb{R}^{T\times H\times W\times C}$ is first partitioned into a sequence of $N_{t}$ non-overlapping temporal clips. Each clip is then sampled into $F$ frames. These frames are further decomposed into a grid of non-overlapping patches, which are linearly projected into patch embeddings. A special ‘[CLS]‘ token is prepended to the sequence. The entire sequence is then processed by a series of Transformer blocks that apply self-attention across both spatial and temporal dimensions. The output of the final Transformer layer corresponding to the ‘[CLS]‘ token for each clip serves as its high-level representation. The collection of these clip representations forms the global context feature set, $Z_{global}$.
@@
[…]

### v2

@@ 1 Introduction
The proliferation of video data has positioned it as a primary medium for information exchange and environmental perception, driving a significant evolution in computer vision research (fu2024survey; li2024multimodal). Historically, the field has achieved remarkable success in discriminative tasks—moving from foundational image recognition (dosovitskiy2021an) to complex action and gesture recognition in videos (abu-farha-2021-time). This progress is mirrored in specialized domains such as sign language translation, where hierarchical models have demonstrated powerful recognition capabilities , and in fine-grained sensing technologies like WiFi-based gesture and activity recognition, which have become increasingly robust and resistant to interference (gu2023attention; zhang2025wiopen). However, this paradigm of "recognition" primarily addresses the question of "what" is happening in a visual scene. The community is now shifting towards a more profound challenge: moving from mere perception to genuine cognition (zellers2019vcr). This involves enabling machines to reason about "why" an event is occurring and to predict "what" might happen next—a task that requires a level of understanding far beyond statistical pattern matching.
@@
The primary obstacle to achieving this cognitive leap is the "knowledge gap." Current models, despite their architectural sophistication, often operate within a closed world, lacking the vast repository of commonsense, physical intuition, and social knowledge that humans effortlessly apply. For instance, while a model might accurately classify a video clip as "a person picking up an egg and flour," it typically fails to infer the underlying intent, such as "the person is about to bake a cake." This limitation becomes particularly salient when considering the demands of next-generation applications. Advanced human-machine systems (zhang2025wiopen), affective computing that aims to suppress label noise for reliability (zhang2025resup), and even emerging fields like psychological understanding via language models all necessitate a deeper, causal understanding of events. Similarly, while we can now capture nuanced physiological data through commodity devices (zhao2025wi), interpreting this data in the context of complex human activities requires advanced reasoning. The core challenge, therefore, is to imbue visual models with this external world knowledge.
@@
To bridge this gap, we propose a novel framework that synergistically fuses the capabilities of two of the most powerful paradigms in modern AI: Vision Foundation Models (VFMs) and Large Language Models (LLMs). Our approach leverages VFMs, such as those built upon the principles of Vision Transformers (dosovitskiy2021an) and trained with multimodal supervision (radford2021learning; wang2022internvideo), to serve as the "eyes" of our system. These models are adept at extracting rich, spatiotemporal features and providing a detailed, pixel-level perception of the visual world. Concurrently, we employ a pre-trained LLM, such as LLaMA (touvron2023llama), as the "brain"—a reasoning core endowed with a vast reservoir of world knowledge, causal relationships, and abstract concepts, as demonstrated in seminal works like GPT-4 (bubeck2023sparks). The crux of our methodology lies in a meticulously designed fusion mechanism that translates the continuous, unstructured visual evidence from the VFM into a discrete, language-like format that the LLM can process and reason upon. This approach is inspired by the success of pioneering vision-language architectures like Flamingo (alayrac2022flamingo) and BLIP-2 (li2023blip), but is explicitly tailored for complex event-level inference and prediction, moving beyond simple captioning or direct question-answering.
@@
This work makes several significant contributions to the advancement of cognitive video understanding. First and foremost, we introduce a novel and effective framework that, for the first time, cohesively integrates a state-of-the-art Vision Foundation Model with a large-scale LLM to perform high-level event reasoning and prediction, moving decisively beyond simple recognition tasks. Central to this framework is our second contribution: the design of a lightweight yet powerful cross-modal fusion module. This component acts as an efficient information bottleneck, aligning rich visual features with the LLM’s semantic space, which not only enables complex inference but also ensures that the model’s reasoning is firmly grounded in direct visual evidence. To validate our approach, our third contribution is a set of extensive experiments conducted on multiple challenging video reasoning benchmarks. The results demonstrate that our model significantly outperforms existing state-of-the-art methods and, notably, exhibits remarkable zero-shot capabilities in predicting future events, underscoring the profound benefits of transferring world knowledge from LLMs. We believe the principles developed herein also hold considerable promise for enhancing related multimodal tasks, such as visual dialogand audio-visual event analysis .
@@ 2.1 Foundations of Visual Representation Learning
The journey towards meaningful video understanding begins with the extraction of powerful visual representations. Early successes were largely dominated by Convolutional Neural Networks (CNNs), which demonstrated exceptional capabilities in hierarchical feature extraction for images. However, the advent of the Transformer architecture, particularly the Vision Transformer (ViT) (dosovitskiy2021an), marked a paradigm shift. By treating an image as a sequence of patches, ViTs enabled the application of self-attention mechanisms to capture global context, a feat that was challenging for CNNs with their limited receptive fields. This architectural innovation laid the groundwork for a new generation of foundation models. For video, this principle was extended to the temporal domain, leading to powerful video foundation models like InternVideo (wang2022internvideo), which learn generalizable representations from massive datasets through a combination of generative and discriminative objectives. The development of such backbones is a field of research in itself, with ongoing efforts to improve efficiency and effectiveness, for example, through advanced model compression techniques like multi-objective convex quantization or by designing specialized architectures for specific tasks like crowd counting.
@@ 2.2 From Recognition to Spatiotemporal Understanding
Building upon these powerful visual backbones, research has moved beyond simple classification towards a more nuanced understanding of spatiotemporal dynamics. This evolution is evident in tasks that require localizing events in both space and time. Video grounding, which aims to find a specific video segment corresponding to a textual query, is a prime example. Recent works like have focused on developing efficient temporal filtering mechanisms to precisely identify these moments. The ambition has further expanded to generating structured summaries of long-form videos, such as creating distinct chapters, a task addressed by large-scale datasets and models like VidChapters-7M (g-t-2023-vidchapters). This fine-grained temporal understanding is fundamental to our work, as reasoning about event causality and prediction necessitates a precise grasp of "when" things happen. A related task, text-to-video retrieval, further underscores the importance of fine-grained alignment, with recent benchmarks like Ground-A-Video (huang2024groundavideo) pushing the state of the art in accurately matching semantic queries to video content.
@@ 2.3 The Rise of Vision-Language Models (VLMs)
The true catalyst for advanced visual reasoning was the effective fusion of vision and language. The development of CLIP (radford2021learning) demonstrated that a shared embedding space for images and text, learned via massive-scale contrastive pre-training, could enable remarkable zero-shot transfer capabilities. This breakthrough paved the way for a host of large-scale Vision-Language Models (VLMs). Early influential models like Flamingo (alayrac2022flamingo) introduced gated cross-attention layers to inject visual features into a pre-trained and frozen language model, showcasing impressive few-shot learning. This "frozen LLM" paradigm was further explored in works like (tsimpoukelli2021multimodal), which highlighted the potential of this frugal approach. Architectures like BLIP-2 (li2023blip) advanced this idea by introducing a lightweight "Q-Former" module to bridge the modality gap between a frozen image encoder and a frozen LLM, proving to be a highly effective and parameter-efficient strategy. The field continues to expand the scope of fusion, aiming to create omni-perception models like VALOR (chen2024valor) and LanguageBind (zhu2023languagebind), which align not just vision and text, but also audio, depth, and thermal data into a unified semantic space. This trend towards multimodal fusion is not limited to mainstream sensors; innovative research has shown the potential of fusing commodity WiFi signals with vision for tasks like emotion recognition (gu2023wife), illustrating a broader principle of synergistic sensing that our work builds upon.
@@ 2.4 Large Language Models for Video Understanding and Reasoning
The confluence of powerful VLMs and the demonstrated reasoning capabilities of LLMs (touvron2023llama; bubeck2023sparks) has given rise to the current frontier: LLM-powered video understanding. A first wave of models, often framed as "video assistants," focused on enabling dialogue about video content. Models such as Video-LLaMA (zhang2023videollama), Video-ChatGPT (maaz2023videochatgpt), and Chat-UniVi (jin2024chatunivi) demonstrated how a video encoder could be connected to an LLM to answer questions, generate descriptions, and hold conversations about a video’s content. LLaViDA (zhao2023llavida) further explored enhancing this understanding via in-context learning.
Subsequently, research has pivoted towards enabling more complex and structured reasoning. SeViLA (yu2023sevila) introduced a self-chained question-answering approach, encouraging the model to break down problems into smaller, manageable steps. This aligns with broader trends in NLP, such as training models along explicit reasoning paths (geng2023ponderbeforespeaking). Perhaps the most innovative approach is ViperGPT (suris2023viperGPT), which empowers an LLM to write and execute Python code that calls upon various vision APIs, effectively transforming the LLM into a cognitive orchestrator that can answer complex visual queries by composing modular tools. As the complexity of reasoning increases, so does the demand for handling longer contexts. Models like LaVi-L (zhan2023la) and the memory-augmented Stammer (li2024stammer) are specifically designed to address the challenges of long-form video understanding, which is crucial for tracing causal chains over extended periods.
This rapid progress has also necessitated a critical examination of model limitations, particularly the issue of "hallucination," where models generate factually incorrect or ungrounded text. Research like Woodpecker (yin2023woodpecker) is now focused on developing methods to detect and correct these hallucinations, a critical step towards building reliable systems. The ultimate ambition is to create unified, any-to-any multimodal models like Emu2 (sun2023emu2), NExT-GPT (wu2023nextgpt), and Google’s Gemini (geminiteam2023gemini), which aim to seamlessly process and generate content across nearly all modalities. This includes extending reasoning into the third dimension, as explored by Chat-3D-v2 (yao2024chat3dv2), and leveraging novel fusion architectures like AMAM’s modality-adaptive mind (lee2024amam). The fusion principles explored in these works can even find analogues in other domains, such as the multi-view, multi-epoch architecture of SUTRA for speech processing (sainath2024sutra), suggesting a universal trend in multimodal AI.
@@ 2.5 Event Prediction and World Models
The "prediction" component of our work is directly related to the long-standing challenge of video forecasting. Traditional approaches often focused on low-level prediction, such as generating future pixels. Diffusion models, as seen in MCVD (voleti2022mcvd), have recently shown great promise in generating high-fidelity future frames. However, our focus is on higher-level semantic prediction. This aligns with research in human trajectory prediction, where models like V-STF (lee2024vstf) learn to anticipate future movements by fusing social and temporal cues.
The most ambitious vision for prediction is embodied by the concept of "World Models." Pioneering work like DreamerV3 (hafner2023mastering) has shown that an agent can learn a robust internal model of its environment’s dynamics, allowing it to "dream" or simulate future outcomes to plan its actions effectively. This represents a shift from reactive prediction to proactive simulation. The recent Genie model (genie2024) takes this a step further, learning to generate entire interactive, playable 2D worlds from a single image. While our work does not build an explicit world model, it shares the same spirit: leveraging accumulated knowledge to make informed predictions about future states. The remarkable capabilities of generative models like VideoPoet (kondratyuk2023videopoet) to synthesize coherent, dynamic video from text further suggest that these models are implicitly learning deep, predictive representations of the world.
@@ 2.6 Applications, Benchmarks, and Broader Context
The ultimate goal of video reasoning and prediction is to enable intelligent applications and systems. A major beneficiary is embodied AI and robotics. The paradigm has shifted from passive video analysis to training active agents that can perceive, reason, and act in the physical world. Landmark models like RT-2 (brohan2023rt2) and the generalist Octo transformer (octo_blog) demonstrate that a single vision-language-action model can be trained to control robots for a variety of tasks. This requires not only understanding instructions but also organizing and planning complex actions, a challenge addressed by agents like LEO (ma2024leo). The importance of external knowledge in these embodied tasks is highlighted by specialized benchmarks like OK-VILA (sarch2023okvila).
Progress in this field is critically dependent on challenging and well-designed benchmarks. Datasets like CLEVRER (yi2020clevrer) specifically target causal and physical reasoning, while Test of Time (momeni2023test) focuses on evaluating temporal understanding. The Ego-Exo4D dataset (gandhi2023egoexo4d) pushes the frontier by providing synchronized first-person and third-person views of the same event, demanding a more holistic, cross-view understanding. While much research relies on traditional visual data, parallel advancements in alternative sensing modalities are creating new opportunities. Technologies using commodity WiFi (gu2023attention) and RFID (feng2025rf) are now capable of fine-grained activity and even keystroke detection (huang2024keystrokesniffer). These rich data streams, often generated in complex real-world settings like robotic vehicle perception (wang2023drive), require the same, if not more, sophisticated reasoning models to be interpreted meaningfully. This is especially true in healthcare, where vision-based assessment of Parkinson’s tremors (wu2024visionbased) or WiFi-based pulmonary function analysis (zhao2025wi) requires a deep understanding of subtle temporal patterns.
Finally, training these massive models on diverse, real-world data presents its own challenges, leading to research in areas like federated learning to handle decentralized data and heterogeneous networks, with frameworks like Finch enabling neural architecture search in such settings (liu2024finch). Our work is situated at the intersection of these advancements, aiming to leverage foundational models and advanced reasoning to create a system that not only understands video but can also anticipate its future, with broad implications for all these application domains.
@@ 3.1.1 Spatiotemporal Feature Extraction
We employ a pre-trained video foundation model, specifically a variant of the Vision Transformer (ViT) architecture adapted for video, such as InternVideo (wang2022internvideo), as our primary spatiotemporal encoder. The input video $V\in\mathbb{R}^{T\times H\times W\times C}$ is first partitioned into a sequence of $N_{t}$ non-overlapping temporal clips. Each clip is then sampled into $F$ frames. These frames are further decomposed into a grid of non-overlapping patches, which are linearly projected into patch embeddings. A special ‘[CLS]‘ token is prepended to the sequence. The entire sequence is then processed by a series of Transformer blocks that apply self-attention across both spatial and temporal dimensions. The output of the final Transformer layer corresponding to the ‘[CLS]‘ token for each clip serves as its high-level representation. The collection of these clip representations forms the global context feature set, $Z_{global}$.
@@
where $V_{clip_{i}}$ is the $i$-th clip of the video, $\text{VideoTransformer}(\cdot)$ denotes the forward pass through the vision backbone, and $D_{v}$ is the dimension of the visual features.